OpenStudy (anonymous):

Use differentials to estimate the amount of tin in a closed tin can with diameter 10 cm and height 15 cm if the tin is 0.04 cm thick. (Round your answer to two decimal places.)

OpenStudy (irishboy123):

try this: the volume of the can is: \(\large V = \frac{\pi D^2}{4} \, h \) so: \(dV = V_D \, dD + V_h \, dh\) \( = \frac{\pi \, D}{2} \, h \, dD + \frac{\pi D^2 }{4} \, dh\) \(\large dD = -\frac{0.04}{2} = -0.02, \, \, dh = -0.04\)